A Mathematics professor is teaching both a morning and an afternoon section of introductory calculus. Let A={the professor gives a bad morning lecture} and B={the professor gives a bad afternoon lecture}. If P(A) = 0.3, P(B) = 0.2 and P(A ∩ B) = 0.1. Calculate the following probabilities:

(a) P(B|A)

(b) P(B'|A)

(c) P(B|A' )

(d) P(B'|A')

(e) If at the conclusion of the afternoon class, the professor is heard to mutter “what a rotten lecture", what is the probability that the morning lecture was also bad?


1
Expert's answer
2022-02-02T04:54:12-0500

Given that,

"p(A) = 0.3, p(B) = 0.2, p(A \cap B) = 0.1".


"a)"

"p(B|A)={P(A\cap B)\over p(A)}={0.1\over0.3}={1\over3}"


"b)"

"p(B'|A)=1-p(B|A)=1-{1\over3}={2\over3}"

"c)"

From part "b" above, we can obtain "p(A\cap B')" as follows,

"p(B'|A)={p(A\cap B')\over p(A)}"

So,

"{2\over3}={p(A\cap B')\over0.3}\implies p(A\cap B')={2\over3}\times0.3=0.2"

Now,

"P(B|A' )={p(A\cap B')\over p(A')}={p(A\cap B')\over1- p(A)}={0.2\over 1-0.3}={0.2\over 0.7}={2\over7}"


"d)"

"P(B'|A')=1-p(B|A')=1-{2\over7}={5\over7}"


"e)"

We determine the conditional probability

"p(A|B)={p(A\cap B)\over p(B)}={0.1\over 0.2}={1\over2}"

Therefore, the probability that the morning lecture is bad given that the afternoon lecture is bad is "{1\over2}."